# PriorDiffusion: Leverage Language Prior in Diffusion Models for Monocular Depth Estimation

### v3

@@ C.2. Ablation for prompts to generate text
As demonstrated in Figure 8, we evaluate PriorDiffusion and Marigold with different denoising steps during inference. The extracted performances are shown in Table 4. With more denoising steps, performance gradually improves. PriorDiffusion consistently outperforms Marigold across various denoising steps, converging in just 10 steps, while Marigold requires 25 steps. This suggests that the language prior could offer a good initialization, speeding up the denoising process and accelerating convergence.
@@
Steps NYUv2 KITTI ETH3D ScanNet $\delta_{1}$ AbsRel↓ $\delta_{1}$ AbsRel↓ $\delta_{1}$ AbsRel↓ $\delta_{1}$ AbsRel↓ Marigold 1 48.8 33.8 25.2 59.1 50.1 37.5 60.3 25.7 2 78.7 16.9 72.2 18.0 86.0 12.7 72.1 19.2 3 92.6 8.2 77.7 15.3 92.0 8.6 87.1 11.4 4 94.2 7.0 80.4 14.1 93.6 7.6 90.7 9.3 5 94.7 6.7 82.4 13.4 94.3 7.3 91.9 8.6 10 95.1 6.4 86.5 11.9 94.9 6.9 93.2 7.8 15 95.1 6.4 87.4 11.6 95.0 6.9 93.4 7.7 20 95.2 6.3 88.0 11.4 95.0 6.9 93.7 7.5 25 95.3 6.3 88.2 11.3 95.0 6.9 93.7 7.5 50 95.3 6.3 88.5 11.2 95.0 6.9 93.8 7.5 PriorDiffusion 1 48.8 33.8 25.2 59.1 50.1 37.5 60.3 25.7 2 83.2 14.1 74.7 16.9 86.5 12.0 75.0 17.8 3 94.3 7.1 81.2 13.9 93.0 7.7 90.6 9.5 4 95.5 6.3 84.7 12.4 94.7 6.9 93.5 7.8 5 95.7 6.0 87.0 11.6 95.3 6.6 91.9 8.6 10 96.0 5.9 89.9 10.5 95.7 6.5 94.9 6.8 15 96.0 5.9 90.1 10.4 95.8 6.5 94.9 6.8 20 96.0 5.9 90.3 10.4 95.8 6.5 94.9 6.7 25 96.0 5.9 90.3 10.4 95.8 6.5 94.9 6.7 50 96.0 5.9 90.3 10.4 95.8 6.5 94.9 6.7
@@
Table 4: Performance for different denoising steps. PriorDiffusion consistently outperforms Marigold across various denoising steps, with a significantly faster convergence speed for the diffusion process.
C.4. Template prompt comparison.
In situations where user-provided descriptions are not available, we aim to explore whether our method can still perform effectively using either a blank input or standardized template prompts. If no specific language description that aligns with the scene is given by the user, we could instead use some pre-defined template prompt, such as “An image” or more elaborate prompts like, “A complex 3D scene with varying objects at different distances.”, as input of PriorDiffusion to maintain its performance. ” These templates, as presented in Table 5, assist in preserving the model’s performance when language input is not practical. Consequently, using template prompts as input during inference may differ from the training domain, leading to a significant covariate shift in language descriptions. This issue can be mitigated by incorporating template prompts as part of the input during training, enabling PriorDiffusion to better adapt to various scenarios where meaningful language descriptions are not feasible. Nevertheless, PriorDiffusion can achieve comparable or even better performance with Marigold when language input is not feasible, thus using a fixed prompt instead, showing that simply training with language can improve the performance of the depth estimator.
@@
Method NYUv2 KITTI ETH3D $\delta_{1}$ AbsRel↓ $\delta_{1}$ AbsRel↓ $\delta_{1}$ AbsRel↓ Blank text input 95.5 6.2 89.3 10.9 95.0 6.9 “An image” 95.7 6.1 89.8 10.7 95.1 6.8 Template A 95.8 6.0 89.9 10.7 95.3 6.8 Template B 95.7 6.1 89.8 10.7 95.2 6.8 Template C 95.8 6.1 89.8 10.7 95.3 6.8 Marigold 95.7 6.1 89.7 10.7 95.4 6.9 With text input 96.0 5.9 90.3 10.4 95.8 6.5
@@
Table 5: Template prompt comparison. Metrics are reported as percentages. PriorDiffusion can achieve comparable or even better performance with Marigold when language input is not feasible, thus using a fixed prompt instead. The last row is the original PriorDiffusion with text input.
C.5. Obtain language description.
To generate text descriptions for images, for training images, we use two different version of LLaVA v1.6 [40], Mistral and Vicuna, each with 5 different prompts, to generate 10 text descriptions for each training image:
@@
This prompt is designed to elicit responses that include essential details, such as the positioning of objects, their interactions, and notable features that may impact depth estimation. Note that all training data we use are synthetic data, and by emphasizing “assuming it’s a real-world image,” we ensure that the descriptions align with the types of inputs and scenarios the model will encounter.

### v4

@@ C.3. Different denoising steps
As demonstrated in Figure 8, we Marigold baseline and Marigold with both training text and inference text, with different denoising steps during inference. The performances are shown in Table 4. With more denoising steps, performance gradually improves. Integrating text consistently outperforms the baseline across various denoising steps, converging in just 10 steps, while the baseline requires 25 steps. This suggests that the language can speed up the denoising process and accelerating convergence.
@@
Steps NYUv2 KITTI ETH3D ScanNet $\delta_{1}$ AbsRel↓ $\delta_{1}$ AbsRel↓ $\delta_{1}$ AbsRel↓ $\delta_{1}$ AbsRel↓ Marigold 1 48.8 33.8 25.2 59.1 50.1 37.5 60.3 25.7 2 78.7 16.9 72.2 18.0 86.0 12.7 72.1 19.2 3 92.6 8.2 77.7 15.3 92.0 8.6 87.1 11.4 4 94.2 7.0 80.4 14.1 93.6 7.6 90.7 9.3 5 94.7 6.7 82.4 13.4 94.3 7.3 91.9 8.6 10 95.1 6.4 86.5 11.9 94.9 6.9 93.2 7.8 15 95.1 6.4 87.4 11.6 95.0 6.9 93.4 7.7 20 95.2 6.3 88.0 11.4 95.0 6.9 93.7 7.5 25 95.3 6.3 88.2 11.3 95.0 6.9 93.7 7.5 50 95.3 6.3 88.5 11.2 95.0 6.9 93.8 7.5 Marigold + Text (Training & Inference) 1 48.8 33.8 25.2 59.1 50.1 37.5 60.3 25.7 2 83.2 14.1 74.7 16.9 86.5 12.0 75.0 17.8 3 94.3 7.1 81.2 13.9 93.0 7.7 90.6 9.5 4 95.5 6.3 84.7 12.4 94.7 6.9 93.5 7.8 5 95.7 6.0 87.0 11.6 95.3 6.6 91.9 8.6 10 96.0 5.9 89.9 10.5 95.7 6.5 94.9 6.8 15 96.0 5.9 90.1 10.4 95.8 6.5 94.9 6.8 20 96.0 5.9 90.3 10.4 95.8 6.5 94.9 6.7 25 96.0 5.9 90.3 10.4 95.8 6.5 94.9 6.7 50 96.0 5.9 90.3 10.4 95.8 6.5 94.9 6.7
@@
Table 4: Performance for different denoising steps. Integrating text consistently outperforms the baseline across various denoising steps, with a significantly faster convergence speed for the diffusion process.
C.4. Template Prompt Comparison
When training with text, we also consider inference scenarios where user-provided descriptions are unavailable. We therefore explore whether the model can still perform comparably by using either a blank input or standardized template prompts. Specifically, we use the Marigold model trained with text and evaluate the effect of several predefined prompts—ranging from simple ones such as blank string “”, simple words like “An image”, to more descriptive ones like “A complex 3D scene with varying objects at different distances.”, as inputs to the diffusion-based depth estimator to maintain its performance.
As presented in Table 5, these template prompts help preserve the model’s performance when explicit language input is not feasible. The results show that the model achieves comparable, or even better, performance than the Marigold baseline when using fixed prompts instead of user-provided text. This finding suggests that, even when user-provided descriptions are unavailable, incorporating language during training itself might enhance the depth estimator’s generalization and overall performance.
@@ C.4. Template Prompt Comparison
Method NYUv2 KITTI ETH3D $\delta_{1}$ AbsRel↓ $\delta_{1}$ AbsRel↓ $\delta_{1}$ AbsRel↓ Blank text input 95.5 6.2 89.3 10.9 95.0 6.9 “An image” 95.7 6.1 89.8 10.7 95.1 6.8 Template A 95.8 6.0 89.9 10.7 95.3 6.8 Template B 95.7 6.1 89.8 10.7 95.2 6.8 Template C 95.8 6.1 89.8 10.7 95.3 6.8 Marigold 95.7 6.1 89.7 10.7 95.4 6.9 With text input 95.9 5.9 90.6 10.4 95.7 6.5
@@
Table 5: Inference using fixed template text input. The results show that the model achieves comparable, or even better, performance than the Marigold baseline when using fixed prompts instead of user-provided text. This finding suggests that, even when user-provided descriptions are unavailable, incorporating language during training itself might enhance the depth estimator’s generalization and overall performance.
C.5. Ablation on the Number of Text Captions per Image
We test Marigold’s performance with different numbers of captions provided for each image, to test whether increasing the number of text provided during training and improve the performance. To generate text descriptions for images, for training images, we use two different version of LLaVA v1.6 [liu2023improved], Mistral and Vicuna, each with 5 different prompts, to generate 10 text descriptions for each training image:
@@ C.5. Ablation on the Number of Text Captions per Image
This prompt encourages generated responses to include essential details for depth estimation, such as the positioning of objects, their interactions, and notable features that may impact depth estimation. Note that all training data we use are synthetic data, and by emphasizing “assuming it’s a real-world image,” we ensure that the descriptions align with the types of inputs and scenarios the model will encounter.
When multiple captions are available for each image, one is randomly selected during training. Shown in Table 6, as the number of captions increases, performance saturates and the improvement is marginal. One possible explanation is that different captions provides similar descriptions of the scene in terms of attributes essential for depth estimation, such as object types, sizes, spatial relationships, and scene structure, since those captions are prompted to cover all those essential details. Generating additional captions does not introduce additional information, thus leading to only marginal improvements in monocular depth estimation accuracy.
Method NYUv2 KITTI ETH3D $\delta_{1}$ AbsRel↓ $\delta_{1}$ AbsRel↓ $\delta_{1}$ AbsRel↓ 1 Caption Per Image 95.9 5.9 90.6 10.4 95.7 6.5 2 Captions Per Image 95.9 5.9 90.5 10.4 95.7 6.5 5 Captions Per Image 96.0 5.9 90.4 10.4 95.8 6.5 10 Captions Per Image 96.0 5.9 90.3 10.4 95.8 6.5
Table 6: Training with different numbers of text captions per image. For images annotated with multiple captions, a caption is randomly sampled for each training iteration. While adding more captions initially slightly improves performance, the benefit quickly saturates, yielding only minor gains beyond a certain point.